OpenStudy (anonymous):

The length of a rectangle is twice the width the area is 242 yd. Find the length and the width.

OpenStudy (anonymous):

length l=2w (w=width), right? and area: A=l*w

OpenStudy (anonymous):

A=2w*w=242 so, w^2=121 or w=11 can you find l (length ) ?

OpenStudy (anonymous):

length l=2w=2*11=22

OpenStudy (anonymous):

Area of rectangle: length * width = 242 yd length = 2 * width length * width = 242 yd (2 * width) * width = 242 yd 2 * width^2 = 242 yd width^2 = 121 yd width =11 yd length = 22 Check: lenght * width = 22 * 11 = 242

OpenStudy (anonymous):

Sharon!!!